# Distribution Matching Distillation Meets Reinforcement Learning

### v3

@@ Abstract
Distribution Matching Distillation (DMD) distills a pre-trained multi-step diffusion model to a few-step one to improve inference efficiency. However, the performance of the latter is often capped by the former. To circumvent this dilemma, we propose DMDR, a novel framework that combines Reinforcement Learning (RL) techniques into the distillation process. We show that for the RL of the few-step generator, the DMD loss itself is a more effective regularization compared to the traditional ones. In turn, RL can help to guide the mode coverage process in DMD more effectively. These allow us to unlock the capacity of the few-step generator by conducting distillation and RL simultaneously. Meanwhile, we design the dynamic distribution guidance and dynamic renoise sampling training strategies to improve the initial distillation process. The experiments demonstrate that DMDR can achieve leading visual quality, prompt coherence among few-step methods, and even exhibit performance that exceeds the multi-step teacher.
@@ 1 Introduction
Diffusion models have achieved unparalleled quality in visual generation tasks [dit, sd3, flux, lumina2, sdxl]. Yet their sampling process typically demands tens of iterative denoising steps, each involving a full forward pass through a large neural network, which makes high-resolution text-to-image synthesis costly both in time and computation.
To accelerate sampling speed, a surge of work [diffinstruct, dmd, lcm, hypersd, dmd2, flashdiff] has sought to distill the original diffusion model into a generator that needs only a handful of steps (e.g, 4). These approaches can be roughly divided into two types: distillation from the trajectory and distillation from the distribution. The former method operates on the instance level, aiming to simulate the original teacher diffusion generation trajectories in as few steps as possible [lcm, piflow]. While the latter works at the distribution level, striving to match the output distribution of the few-step student generator to the original multi-step one’s [diffinstruct, dmd].
In this study, we focus on distribution matching distillation (DMD) as it has been verified in large-scale settings [dmd2, dmdx] and has already been used in some leading products (e.g, Seedream 4.0 [seedream4]). However, one notable thing is that in DMD, the student model aims to match the distribution of the teacher model, which also means that the ability of the student model is often limited by the teacher. To remedy this, most works (e.g., DMD2 [dmd2]) introduce an additional GAN branch for guidance [gan]. However, training with GAN not only renders the training process dependent on external high-quality visual data, but also introduces training instability inherent in GAN [wgan, improvedgan].
To this end, we propose DMDR, which combines distribution matching distillation (DMD) with recent powerful diffusion reinforcement learning (RL) [refl, flowgrpo, diffusiondpo, ddpo]. In a nutshell, DMDR enables the student model to obtain supervision signals beyond the teacher model without relying on external image data, and makes it possible for the student to surpass the teacher. Contrary to the prevailing belief that RL should be conducted after obtaining the few-step generator, we perform RL concurrently with the distillation process. We show that such a combination is mutually beneficial: (1) As for DMD, RL can help to reshape some existing low-probability modes by guiding them toward higher reward regions, ensuring these modes are covered during the distillation process, thereby mitigating the risk of zero forcing. (2) As for RL, DMD loss itself can serve as a more effective regularization than its traditional counterparts, thereby mitigating the risk of reward hacking.
Meanwhile, since obtaining valid reward scores requires the few-step model to possess a basic generation capability already, we design two dynamic training strategies to facilitate a faster and better “cold start” for DMDR: (1) DynaDG, which strengthens the initial distributional overlap between the teacher and student model; (2) DynaRS, which enhances the student’s global modeling of the teacher’s distribution during the early training phase.
@@
Taken together, our experimental results show that our method leads to state-of-the-art few-step generative models that can outperform their teacher, using as few as 4 sampling steps. Moreover, our method is not only compatible with different models (e.g., flow-based, denoising-based) but also with various RL algorithms (e.g., ReFL, DPO, GRPO). This ensures the long-term effectiveness of our method as the multi-step model and RL algorithm evolve.
@@
  • We design a dynamic cold start stage for faster and better distillation in the initial phase.
  • We validate our method on various models and RL algorithms with all of which achieve excellent performance.
Figure 1: More Images generated by Z-Image-Turbo [zimage] distilled through our DMDR.
2 Related Work
We highlight key related studies here, and leave the discussions of others in Appendix A.
Distribution Matching Distillation. Distribution Matching Distillation (DMD) [dmd] is the first work to successfully apply the principle score-based distillation [score-base-diff, prolificdreamer, dreamfusion] to large-scale text-to-image models. Intuitively, it strives to ensure that any sample realized by the student at a given noise level occurs with exactly the same probability as it would under the teacher’s distribution, thereby preserving the multi-step generative priors in the few-step model. From then on, a lot of follow-up work emerged [dmd2, dmdx, tdm, sd3.5flash, f-distill], for example, DMD2 [dmd2] employs a discriminator to align the student model distribution with a specific target distribution like GAN [gan] does; TDM [tdm] incorporates DMD loss in the sampling process of the student model for better alignment; ff–distill [f-distill] replaces the original reverse Kullback–Leibler (KL) divergence to the proposed $f$-divergence for covering different divergences with different properties. Our work also starts with DMD, but we don’t focus on how to better “imitate” the teacher. Instead, we aim to incorporate reinforcement learning in the distillation process to enable the student to outperform the teacher.
Reinforce Learning for Diffusion Models. Reinforce learning helps to align diffusion models to human preferences by training a reward model and using it to guide generation [ddpo, flowgrpo, refl, srpo, diffusiondpo, dancegrpo]. There are many algorithms to conduct RL, for example, DDPO [ddpo] adapts PPO via image log-likelihoods; ReFL [refl] bypasses likelihoods by optimizing outputs with frozen-reward gradients; Diffusion-DPO [diffusiondpo] adapts DPO to diffusion for paired human preference data; and recent GRPO extensions [flowgrpo, dancegrpo] use GRPO to diffusion models by coupling the training loss with SDE samplers. Although significant progress has been made in RL for diffusion models, most of the work has focused on multi-step models. Here, we find that these algorithms are also adapted to the few-step model when carried out in conjunction with distribution matching distillation.
Figure 2: Overview of DMDR, which contains three key elements: (1) A DMD framework to optimize the generator by utilizing the gradient derived from an implicit distribution matching objective; (2) A RL branch to concurrently incorporate reward feedback from the reward model; (3) Two dynamic training strategies to facilitate more efficient and effective distillation during the initial phase.
@@ 3 Method
The pipeline of DMDR is shown in Figure 2, which contains three key elements: A DMD branch to optimize the generator using the gradient of an implicit distribution matching objective (Section 3.1); A RL branch to synchronously introduce the reward feedback from the reward model (Section 3.2); Two dynamic training strategies to achieve faster and better distillation in the initial phase (Section 3.3).
@@ 3.1 Preliminary
As our goal is to obtain a few-step model with performance beyond its multi-step teacher, and our method is built upon DMD. Here we provide a brief introduction to DMD:
Distribution Matching Distillation (DMD) [dmd] compresses a multi-step diffusion model (teacher) into a few-step generator (student) $G$ by minimizing the time-averaged approximate Kullback-Leibler (KL) divergence between the real distribution $p_{\text{real},t}$ and the synthetic distribution $p_{\text{fake},t}$. Note that $G$ is optimized via gradient descent, and this gradient admits a compact expression as the difference of two score functions:
@@
3.2 DMDR: Brings RL into DMD
It is worth noting that in DMD, the goal of the student generator is to align the output distribution as closely as possible with the distribution learned by the multi-step teacher. However, this also causes the ability of the student model to be capped by the teacher. Unlike the previous methods [dmd2, magic141] that introduce external real data to remedy this, we address the problem in an image-free manner by introducing the recently popular diffusion RL [refl, diffusiondpo, flowgrpo] in the DMD pipeline. Below, we explain how DMD and RL benefit from each other when training together.
RL Unlocks the Performance of DMD. One limitation of DMD is that the student model’s performance is inherently capped by the teacher model. By incorporating an RL objective, we allow the student model to receive supervision signals beyond merely mimicking the teacher, thus making it possible to surpass the teacher in an image-free training setup.
Figure 3: Illustration for mode seeking process.
Specifically, we integrate the reward models which provide reward signals indicating the specific characteristic (e.g, quality, text-alignment) of a generated image. During the joint training process, the RL component guides the student generator $G$ to produce samples that maximize these rewards. This has a direct impact on the DMD process: (1). Helping to jump out of the bad modes: The teacher’s distribution might contain modes that are less desirable or have lower perceived quality according to the reward model. RL actively discourages the student from generating samples within these low-reward regions, and “pulls” the student’s distribution towards regions in the image space that yield higher reward scores even if they are present in the teacher’s distribution. (2) Helping to mitigate zero forcing: In DMD, there’s a risk of “zero forcing,” where the student might fail to cover certain modes present in the teacher’s distribution, as we show the first row of Figure 3. By actively guiding the student towards high-reward modes, RL helps ensure that these important and preferred modes are reshaped, thus adequately covered during distillation, preventing their accidental omission as shown in the second row of Figure 3.
DMD Can Effectively Regularize RL. In addition to maximizing rewards, another important point in RL is the regularization, which helps to avoid rapid overfitting of the reward (reward hacking). Although this regularization takes on different forms in various methods (e.g, pretrain loss in ReFL [refl], KL divergence loss in Flow-GRPO [flowgrpo]), essentially, they all aim to let the training model be close to the reference model (always the one before RL). In the multi-step model, this is tractable because there is a pre-trained reference model. But in the few-step model, the situation is somewhat different. As shown in Figure 4, when RL is naively performed after distillation, the goal of regularization becomes to close the new model’s distribution $P_{\text{fake-new}}$ to the distilled few-step model’s $P_{\text{fake-ref}}$. However, as mentioned above, some modes are inevitably neglected after being distilled, thus $P_{\text{fake-ref}}$ cannot serve as an effective regular expression, as discovered by other works [hypersd] that these lead to quick hacking. However, if RL and DMD are carried out simultaneously, this problem will be easily solved. The DMD loss (Equation 1) continuously pulls the student generator’s distribution ($P_{\text{fake}}$) towards the multi-step teacher’s distribution ($P_{\text{real}}$). If viewed from the RL side, this also means that at every step of the joint training, the student is being regularized by the comprehensive and robust $P_{\text{real}}$ itself (the solid arrow “DMD + RL” in Figure 4), which reduces error accumulation in regularization and the risk of the reward hacking (see Figure 9).
Putting together, the final loss function can be written as:
$\mathcal{L}=\mathcal{L}_{\text{dmd}}+\mathcal{L}_{\text{rl}},$ (2)
@@ 3.2 DMDR: Brings RL into DMD
where $\mathcal{L}_{\text{rl}}$ is the plug-and-play loss from the RL branch, the gradient can be differentiable type like ReFL, or policy type like GRPO, etc. In practice, we will multiply a coefficient if the two losses are not on the same scale.
3.3 A Dynamic Cold Start Stage for DMDR
While DMDR enables the student to surpass its teacher, the reward model often requires the generator to possess a basic generation capability to produce valid rewards [refl, srpo]. In practice, we observe that reward signals are noisy and unreliable in the early training phase. To this end, we conduct a cold start phase and design two dynamic training strategies based on vanilla DMD to accelerate and improve it.
Figure 4: Visualization of distribution matching directions.
Figure 5: Visual comparison between the teachers, selected competing methods [dmd2, hypersd, ladd], and ours. All images are generated using identical noise. Our model produces images with superior quality and prompt coherence. More comparisons are available in Appendix B.
Dynamic Distribution Guidance (DynaDG). In the initial stage of training, due to the significant disparity in image quality (the generator is unable to generate qualified images), the distribution of these generated samples ($P_{fake}$) is completely different from that of the training samples of the real score estimator ($P_{real}$), which finally leads to the inaccuracies in the KL divergence approximation as we show in Figure 6 (left). This substantial divergence comes from sparse or non-existent overlap between $P_{real}$ and $P_{fake}$, rendering the gradient $\nabla\log P_{real}-\nabla\log P_{fake}$ unreliable for effective optimization. Inspired by a recent study [magic141], we introduce a LoRA set [lora] and inject it not only into the fake score estimator to learn the rapidly evolving distribution of the few-step generator but also strategically into the real score estimator. The core idea behind our DynaDG is to dynamically adjust the distribution estimated in the real score branch. By injecting LoRA into the real score estimator with a smaller scale, we can effectively “pull” the perceived $P_{real}$ towards the nascent $P_{fake}$ distribution, creating a region of “more overlap” as depicted in Figure 6 (right). This increased overlap ensures that the estimated score difference, $\nabla\log P_{real}-\nabla\log P_{fake}$, consistently provides a “right direction estimation and optimization” signal, even when the distributions are initially far apart. Crucially, this LoRA scale for the real score estimator gradually weakens as training progresses. This mechanism allows the real score estimator to progressively revert to accurately representing the true $P_{real}$ as the generator improves and $P_{fake}$ naturally converges towards $P_{real}$.
Figure 6: Illustration for our Dynamic Distribution Guidance.
Dynamic Renoise Sampling (DynaRS). Another important operation in DMD is renoising, which renoises the samples generated by the generator for estimating the score (i.e., the distribution) associated with the specific noise level. The core idea of our DynaRS is to initially bias the sampling of renoise levels $t$ in $F_{t}$ towards higher values, and then gradually transition to a uniform sampling like DMD dose. This has two advantages for a faster and better cold start. First, when biasing to high noise, the forward operator makes noise largely occupy the image, so even bad samples produced by the generator in the initial stage lie in a similar regime for both real and fake estimators, thus improving score reliability and overlap. Second, the estimated score at a higher noise level is more about global structure [lsast, prospect], which can guide the generator to first capture global structure. As the capacity of the generator improves, the renoise bias is progressively reduced, allowing the model to learn finer details by sampling from a wider range of noise levels.
3.4 Distinction from Prior Few-step RL Works
Some prior works also try to introduce RL for the few-step diffusion model [hypersd, ladd, hyperbagel, hypernoise], but they mainly perform it after the distillation, and the problem mentioned by Hyper-SD [hypersd] of quickly deviating from the original domain still occurs. In this work, we illustrate the rationale for conducting the two simultaneously and how DMD loss can effectively regularize RL to mitigate the above problems. Meanwhile, although pioneer work like Diff-Instruct ++ [diffinstruct++] also introduces RL in the distillation process, our motivation and analysis are distinct, with more intuitive illustrations of the mutual benefits of the two, the verification of different diffusion models and RL algorithms at a large scale, as well as the dynamic cold start stage before incorporating RL.

### v4

@@ Abstract
Distribution Matching Distillation (DMD) facilitates efficient inference by distilling multi-step diffusion models into few-step variants. Concurrently, Reinforcement Learning (RL) has emerged as a vital tool for aligning generative models with human preferences. While both represent critical post-training stages for large-scale diffusion models, existing studies typically treat them as independent, sequential processes, leaving a systematic framework for their unification largely unexplored. In this work, we demonstrate that jointly optimizing these two objectives yields mutual benefits: RL enables more preference-aware and controllable distillation rather than uniformly compressing the full data distribution, while DMD serves as an effective regularizer to mitigate reward hacking during RL training. Building on these insights, we propose DMDR, a unified framework that incorporates Reward-Tilted Distribution Matching optimization alongside two dynamic distillation training strategies in the initial stage, followed by the joint DMD and RL optimization in the second stage. Extensive experiments demonstrate that DMDR achieves state-of-the-art visual quality and prompt adherence among few-step generation methods, even surpassing the performance of its multi-step teacher model.
@@ 1 Introduction
Diffusion models pre-trained on large-scale datasets have achieved remarkable performance in visual generation tasks [dit, sd3, flux, zimage, sdxl]. However, their sampling process typically requires numerous iterative denoising steps [ddim, ddpm, flow-matching], each involving a full forward pass through a large neural network. This requirement makes high-resolution text-to-image synthesis computationally demanding. Furthermore, base models often fail to produce images that align closely with human preferences and aesthetic criteria [pickscore, hpsv2, hpsv3].
To enhance practical utility, researchers typically employ two post-training stages. First, step-distillation [diffinstruct, dmd, lcm, hypersd, dmd2, flashdiff], which compresses the original model into a generator capable of few-step sampling (e.g., 4 steps) for efficient sampling. Among such methods, Distribution Matching Distillation (DMD) [dmd] is widely recognized for its effectiveness in large-scale scenarios [dmd2, dmdx, twinflow, senseflow] and has been adopted in prominent industrial applications [seedream4, zimage, flux-2]. Second, reinforcement learning (RL) [refl, flow-dpo, flowgrpo, ddpo, dancegrpo, srpo, diffusionnft, diffusiondpo] aligns the model with human preferences to improve aesthetic quality. Despite progress in both domains, most existing studies address these objectives in isolation. Early efforts to incorporate RL into the distillation process were largely confined to a two-stage paradigm [diffinstruct++, hypersd, pso, hypernoise] (exploring how to conduct RL on an already distilled model). Consequently, there remains a critical gap in the field: a systematic framework to explore the unification and potential synergy between these two “temporarily separated” stages.
To bridge this gap, we propose DMDR, a unified framework that seamlessly integrates DMD and RL into a synergistic training pipeline. Our core insight is that concurrent optimization yields mutual benefits: (1) Preference-Aware Distillation: RL steers the distillation process toward high-reward regions, breaking the ”performance ceiling” inherent in purely mimicking a teacher; (2) Distributional Regularization: The DMD objective serves as a robust regularizer for RL, mitigating reward hacking by maintaining proximity to the teacher’s manifold. Specifically, DMDR operates in two stages. In the initial stage, we implement Reward-Tilted Distribution Matching (RT-DM), which distills toward a reward-tilted teacher distribution to prioritize human-preferred regions. Concurrently, we introduce Dynamic Distribution Guidance (DynaDG) and Dynamic Renoise Sampling (DynaRS) to stabilize the ”cold-start” phase through annealed distributional overlap maximization. In the second stage, we transition to a joint optimization paradigm. In this phase, the RL objective drives aggressive preference alignment, while the DMD mechanism serves as a robust regularizer to prevent reward hacking.
@@
Our experimental results show that our method leads to state-of-the-art few-step generative models. Moreover, our method is not only compatible with different models (e.g., flow-based, denoising-based) but also with various RL algorithms (e.g., ReFL, DPO, GRPO). This ensures the long-term effectiveness of our method as the multi-step model and RL algorithm evolve.
@@
  • We design a Reward-Tilted Distribution Matching and two dynamic distillation training strategies for better integrating RL and distillation in the initial phase.
  • We validate our method on various models and RL algorithms, all of which achieve excellent performance.
@@ 2.1 Preliminary
Distribution matching distillation. DMD [dmd] compresses a multi-step diffusion model (teacher) into a few-step generator (student) $G$ by minimizing the time-averaged approximate Kullback-Leibler (KL) divergence between the real distribution $p_{\text{real},t}$ and the synthetic distribution $p_{\text{fake},t}$. Note that $G$ is optimized via gradient descent, and this gradient admits a compact expression as the difference of two score functions:
@@
Reinforce learning for diffusion models. Although there are various RL algorithms [refl, flowgrpo, diffusiondpo], All their goal is to optimize the diffusion model to generate the samples that can maximize a score given by reward models [hpsv2, pickscore, clipscore]. This score can be expressed as aesthetic, texture, prompt following, and so on. At the same time, a regularization term is often added to balance the trade-off between reward maximization and the deviation from the original model [rlhf]. Thus, the loss can be formulated as follows:
$\mathcal{L}_{\text{rl}}=\mathbb{E}\left[r(\bm{x}_{0})-\operatorname{KL}(p_{new}\|p_{\text{ref}})\right],$ (2)
where $r$ is the reward function, $p_{new}$ and $p_{ref}$ is the distribution of the trainable model and the reference (this can be a pre-train data distribution [refl] or the original model’s [flowgrpo, diffusiondpo, diffusionnft]), $\bm{x}_{0}$ is the clean image that is generated by the training generation model.
2.2 Problem Formulation and Pipeline Overview.
It is noted that DMD effectively reduces the sampling steps of the diffusion model, while RL enables the diffusion model to generate instruction-aligned and human-preferred images. Both of these are crucial for the practical application of the diffusion model. However, at present, the majority of the work [dmdx, dmd2, magicdistillation, flowgrpo, diffusionnft, dancegrpo] only focuses on studying and improving one aspect, without exploring the synergy between the two. And only a small portion of the works [pso, flash-dmd, diffinstruct++, hypersd, hypernoise] attempt to introduce RL into the distillation process, but they still focus on how to conduct RL on a distilled model. This raises a question: Can these two be combined into one stage and benefit each other?
We address this challenge question by proposing DMDR, as shown in Figure 1, the training process of DMDR is divided into two operational stages: Stage 1: RT-DM with Dynamic Distillation. (Section 2.4 and Section 2.5) Stage 2: Joint RL + DMD Optimization(Section 2.3). The following sections will provide the detailed analysis and illustration.
Figure 1: Overview of DMDR. It follows a two-stage training paradigm: Stage 1 (Reward-Tilted Distribution Matching) employs reward-weighted distillation to integrate preference signals into the early distillation phase, stabilized by Dynamic Distribution Guidance (DynaDG) and Dynamic Renoise Sampling (DynaRS). Stage 2 (Joint RL + DMD) performs direct reward maximization regularized by the DMD loss.
2.3 Bring DMD and RL Together with Benefit
We now show formally and empirically that DMD operates as a superior regularization mechanism that subsumes the traditional KL penalty in RL for few-step model, yielding robust joint optimization.
Figure 2: Visualization of the difference of optimization directions.
From the perspective of RL, standard preference optimization requires a penalty (e.g., $\operatorname{KL}(p_{new}\|p_{\text{ref}})$) to prevent reward hacking [refl, flowgrpo]. In few-step models, sequentially applying RL after distillation uses the already-distilled distribution $P_{\text{fake-ref}}$ as the regularizer (Fig. 2). Because distillation inherently loses some modes, $P_{\text{fake-ref}}$ acts as an impoverished anchor, leading to rapid overfitting and mode collapse [hypersd]. More fundamentally, instead of viewing DMD and RL as two disparate losses added heuristically, we reframe the joint objective as maximizing reward under a distribution-level structural constraint. As the joint gradient can be decomposed into two orthogonal objectives [rlhf, score-base-diff, ppo]: a ”reward ascent” direction that increases human preference, and a ”manifold projection” term that corrects off-manifold displacement by pulling the student back toward the teacher’s regions.The DMD gradient $\nabla_{\!\theta}\mathcal{L}_{\text{dmd}}$ corresponds exactly to the maximum likelihood update projecting the student output distribution toward the learned data distribution of the multi-step teacher $P_{\text{real}}$. When combined linearly with the reward gradient $\nabla_{\!\theta}\mathcal{L}_{\text{rl}}$, any displacement induced by $r$ that exits the teacher’s high-density manifold generates a large discrepancy $s_{\text{real}}-s_{\text{fake}}$. The $\mathcal{L}_{\text{dmd}}$ gradient thus projects the reward ascent strictly along the teacher’s support.
Simultaneously, from the perspective of distillation, pure DMD forces the student to uniformly mimic the teacher, implicitly establishing a ”performance ceiling” capped by distilling $P_{\text{real}}$ equally. Incorporating the RL objective addresses this fundamental limitation: RL steers the distillation toward high-reward regions, allowing the student to surpass the teacher while the DMD regularizer ensures semantic integrity is preserved.
Figure 3: Verification of the joint optimization synergy. Directly conduct RL (red) on the distilled model exhibits high training rewards but suffers from reward hacking and semantic collapse (low DPG score). In contrast, the joint DMD+RL paradigm (green) effectively boosts the performance (better than the DMD baseline (blue) in all metric) while preventing reward hacking. Better to zoom in to check the effect.
To validate the analysis, we conduct a comparative study using SD3-M [sd3] as the base model, with rewards provided by HPS v2.1 [hpsv2] and ReFL-based RL [refl] for training. We compare three paradigms: (i) DMD Only (Blue line); (ii) DMD then RL (Sequential: distillation followed by RL in red line); and (iii) DMD then DMD+RL (Joint approach: distillation followed by joint optimization in green line). We evaluate these models across training/test reward curves and the DPG_Bench score [dpg]. As shown in Fig. 3(a), when RL is introduced (after 1000 iterations), the sequential approach achieves the highest training reward. However, this gain is obtained through hacking score. Fig. 3(b) and (c) reveal that this it suffers from severe reward hacking and catastrophic forgetting. While its training reward climbs, its test reward plateaus, and its DPG score collapses. This confirms that without the strong distributional anchor of the teacher, the few-step model quickly drifts into narrow, high-reward modes that sacrifice semantic integrity [hypersd]. In contrast, the joint optimization demonstrates a clear synergistic effect. First, it effectively breaks the ”performance ceiling” of the teacher; the test reward significantly surpasses the DMD-only baseline. Second, by treating the DMD loss as a persistent structural regularizer, the model keeps improving DPG score while increasing aesthetic quality. This indicates that the DMD objective prevents the RL process from hacking, while the RL objective guides the distillation toward more human-preferred regions.
Table 1: Comparison of different training strategies. The joint approach avoids the ”distillation gap” of ”RL then distillation” pipeline, achieving the highest preference score with $\sim 10\times$ rollout speedup.
Method HPS DPG Time
SD3-M, resolution $\times1024$, local batchsize 16 on H100
Multi-Step RL 31.06 86.43 9.24s $\sim$ 10.57s
Multi-Step RL then Distillation 30.42 85.92 -
Few-Step Joint RL and DMD 31.42 86.08 0.43s $\sim$ 1.21s
Based on the empirical synergy observed above, we formally define the optimization objective. The total loss is formulated as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{dmd}}+\lambda_{\text{rl}}\mathcal{L}_{\text{rl}},$ (3)
@@ 2.3 Bring DMD and RL Together with Benefit
where $\mathcal{L}_{\text{dmd}}$ is the original distribution matching loss with the gradient formulated in Eq. 1, and $\mathcal{L}_{\text{rl}}$ is the plug-and-play reward maximization loss from the RL branch (Eq. 1), the gradient can be differentiable type like ReFL, or policy type like GRPO, etc. The hyperparameter $\lambda_{\text{rl}}$ is a balancing coefficient used to trade off the two losses.
2.4 Reward-Tilted Distribution Matching for Integrating RL into the Early Phase of DMD
While we have demonstrated that DMD and RL can be trained jointly with mutual benefit, we still rely on a two-stage paradigm that incorporates RL only after an initial ”cold start” distillation phase [hypersd, diffinstruct++]. This requirement stems from the observation that reward models typically necessitate a baseline generation capability to produce reliable feedback [refl, srpo]. Specifically, in DMD frameworks with flow matching model, the image used for gradient computation is defined as $\bm{x}_{t}-(0-t)G_{\theta}(\bm{x}_{t})$, where $\bm{x}_{t}$ is iteratively sampled. Notably, t can be any training time-step; thus, when training a 4-step model, the clean image can be generated in only one to four steps. In the early training phase, however, such few-step sampling often fails to produce coherent images, rendering reward signals noisy and unreliable for direct optimization.
Figure 4: Effectiveness of RT-DM and two-stage training strategy. In Stage 1, the RT-DM paradigm (red/green) significantly accelerates preference alignment compared to the DMD baseline (blue). In Stage 2, while ”RT-DM only” (red) eventually saturates due to the constraints of the teacher’s manifold, the transition to joint DMD+RL (green) successfully breaks this performance plateau by directly optimizing the reward score. This complete two-stage approach achieves the highest test rewards and DPG scores. Better to zoom in to check the effect.
To address this, we propose a Reward-Tilted Distribution Matching (RT-DM) paradigm to ”softly” integrate reward signals into the distillation process. Specifically, we utilize the reward score as a weighting factor for the DMD loss. This formulation enables a more flexible optimization objective compared to direct reward maximization. This is because the reward score serves as a static weighting factor, we avoid the necessity of backpropagating through the multi-step denoising process, which allows us to evaluate the reward on the complete trajectory of the few-step model (e.g., 4 steps) without the computational overhead of a large gradient graph, bypassing the need for single-step $\bm{x}_{0}$ approximations to make the reward score more precise and highly valuable as a reference for optimization efforts. More profoundly, we are changing which distribution the student learns by smoothly integrating reward signals to make the student not learn toward the vanilla $P_{\text{real}}$, but toward a reward-tilted teacher distribution defined proportionally to $r(x_{0})p_{\text{real}}$. The gradient is formulated as:
$\nabla_{\!\theta}\,\mathcal{L}_{\text{RT-DM}}$ $=\mathbb{E}_{t}\!\left[\nabla_{\!\theta}\operatorname{KL}(p_{\text{fake},t}\|(p_{\text{real},t})r(\bm{x}_{0}))\right]$ (4)
$=-\mathbb{E}_{t}\!\left[\int\!e^{\frac{r(\bm{x}_{0})}{\left\|r(\bm{x}_{0})\right\|}}\Bigl(s_{\text{real}}\bigl(F_{t})\bigr)-s_{\text{fake}}\bigl(F_{t}\bigr)\Bigr)\frac{\mathrm{d}G_{\theta}(z)}{\mathrm{d}\theta}\,\mathrm{d}z\right]\!.$
To evaluate the effectiveness of the proposed soft integration, we compare three configurations in Fig. 4: (i) DMD Only (Blue); (ii) RT-DM for the entire duration (Red); and (iii) RT-DM followed by joint DMD+RL (Green). As shown in Fig. 4(a) and (b), the RT-DM paradigm (Red/Green) provides a significant ”warm start” effect, accelerating preference alignment in the early phase (Stage 1) compared to the DMD-only baseline. This confirms that weighting the DMD loss with reward scores successfully incorporates preference signals even when few-step generations are still maturing, avoiding the noise instability of direct RL. However, the performance of the ”RT-DM only” approach (Red) eventually saturates. We hypothesize that since RT-DM primarily acts by re-weighting the teacher’s distribution gradients, its ability to drive the model toward high-reward regions is inherently constrained by the teacher’s manifold. To further unlock the model’s potential, we introduce a more direct reward maximization phase (Green) in Stage 2. By transitioning to joint DMD+RL optimization once the model has reached a stable baseline, we can leverage direct RL gradients to aggressively push the performance beyond the initial plateau. As a result, the joint approach achieves the highest test rewards and DPG scores, demonstrating that RT-DM serves as an ideal bridge to transition from pure distillation to direct preference alignment.
2.5 Annealed Distributional Overlap Maximization for Better Initial Distillation
During the initial phase of distillation, DMD encounters a classic ”cold start” challenge: the synthetic samples generated by the few-step student exhibit substantial divergence from the real-world pre-training samples. This profound disjointness between $P_{\text{fake}}$ and $P_{\text{real}}$ impedes the teacher model’s ability to estimate reliable scores, rendering the gradient $\nabla\log P_{\text{real}}-\nabla\log P_{\text{fake}}$ unreliable for optimization. To mitigate this, we introduce a unified principle of annealed distributional overlap maximization. We propose two complementary techniques to artificially increase the overlap between $P_{\text{fake}}$ and $P_{\text{real}}$ early in training, progressively relaxing them as the model converges.
Dynamic Distribution Guidance (DynaDG). DynaDG maximizes overlap primarily by actively shifting the target distribution. The poor initial Image quality causes $P_{\text{fake}}$ to be disjoint from $P_{\text{real}}$ (Figure 6, left). To solve this, we inject trainable LoRA modules [lora] into the real score estimator to ”pull” the perceived $P_{\text{real}}$ toward the nascent $P_{\text{fake}}$ (Figure 6, right). This ensures a reliable optimization signal by bridging the manifold gap. As training progresses, we anneal the LoRA scale in the teacher model toward zero, allowing the real score estimator to seamlessly revert to the true $P_{\text{real}}$ once $P_{\text{fake}}$ is safely anchored.
Figure 5: Impact of dynamic training strategies. (a) Qualitative comparisons show that our dynamic strategies facilitate faster global structure building (highlighted in red boxes). (b) Quantitative DPG scores further confirm that DynaDG and DynaRS provide more reliable optimization signals, consistently outperforming the baseline throughout Stage 1. Better to zoom in to check the effect.
Dynamic Renoise Sampling (DynaRS). While DynaDG moves the distributions, DynaRS maximizes overlap by exploiting regions where distributions naturally intersect. By heavily biasing the sampling of renoise levels $t$ toward higher noise values at the start of training, we ensure that both initial samples and reference samples are dominated by Gaussian noise, placing them in a similar regime. At these high noise levels, the score estimator provides highly reliable gradients that emphasize global structure [lsast, prospect]. As the student generator improves, the renoise bias is progressively annealed to uniform sampling, allowing the model to capture fine-grained textural details from a wider range of noise levels.
Figure 6: Illustration for Dynamic Distribution Guidance.
As shown in Figure 5 (b), our dynamic training strategies significantly enhance the initial training phase. Qualitative results presented in Figure 5 left further corroborate these findings, showing a faster global structure building when DynaDG and DynaRS are employed.